please can someone help me. Thank You so much.
Solve -6(3x – 1) = 2(3 – 7x) – 4x. Please show all of your work.

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

-6x+6=6-14x-4x
12x=0
x=0

Answer by Alan3354(69443)

-6(3x – 1) = 2(3 – 7x) – 4x
-18x + 6 = 6 - 14x - 4x
-18x = -18x
No unique solution, x can be any number.